OpenStudy (anonymous):

how to solve 41=2/5x-7?

OpenStudy (lgbasallote):

wait.. is 5x-7 both under 2? or are they separate? if they're separated then isolate x first 2/5x = 41 + 7 2/5x = 48 cross multiply 2x = 240 divide by 2 x = 120

OpenStudy (anonymous):

yes it is the answer I knew the answer but I didn't know how to explain it thanks.